OpenStudy (anonymous):

the formula can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length l in ft how far will the car skid at 70mph ?

OpenStudy (anonymous):

\[2\sqrt{5l}\]

OpenStudy (anonymous):

that is the formula

OpenStudy (anonymous):

\[(70)^2=(2\sqrt{5l})^2\]\[4.900=4*5l\]\[20l=4.900\]\[l=\frac{4.900}{20}\]\[l=245\]